ABCD is a trapezium with ABIIDC.The trapezium has an incircle having radius 6 AD=13,BC=15
FIND 1/4 area ABCD

Grade:10

1 Answers

bharat bajaj IIT Delhi
10 years ago
The diameter of the circle = Perpendicular from BC to AD = 12 = h
Area of Trapezium = 1/2 . h . (AD + BC) = 12.28/2 = 28.6
1/4 area ABCD = 42 units^2
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty